OpenStudy (anonymous):

MEDAL AND FAN FOR ANSWER !! PLEASE HELP !! Admission to the zoo is $6 for children and $9 for adults. On a certain day, 1480 people go to the zoo and $10,500 is collected. How many children and how many adults went to the zoo? children: adults:

OpenStudy (kamibug):

6c + 9a = 10,500 c + a = 1480 c = 1480 - a 6(1480 - a) + 9a = 10,500 8880 - 6a + 9a = 10,500 8880 +3a = 10,500 3a = 1620 a = 540 6c + 9(540) = 10,500 6c + 4860 = 10,500 6c = 5640 c = 940 Children: 940 Adults: 540

OpenStudy (anonymous):

what about this one ? A collection of quarters and nickels is worth $13.70. There are 106 coins in all. How many of each coin are there? Quarters: nickels: @KamiBug

OpenStudy (kamibug):

0.25q + 0.05n = 13.70 q + n = 106 q = 106 - n 0.25(106 - n) + 0.05n = 13.70 Solve this equation ^ for the number of nickels, then substitute in the number you get to the first equation at the top for "n" and solve for "q" the number of quarters.:) Got it from here? I gotta go, sorry.

OpenStudy (anonymous):

can I please have the answer real fast ?